The area os a rectangular feild is 2275m^2. The field is enclosed by 200 m of fence. What are the dimensions of the field?
Answer by checkley77(12844)

2X+2Y=200 SUBTRACT 2Y
2X=200-2Y DIVIDE BY 2
X=100-Y
XY=2,275 REPLACE X BY (100-Y)
(100-Y)Y=2,275
100Y-Y^2-2,275=0
Y^2-100Y+2,275=0
(y-65)(Y-35)=0
Y-65=0
Y=65 ANS.
Y-35=0
Y=35 ANS.